OpenStudy (anonymous):

6x ÷ 3y - 4, if x = 20 and y = 2.

OpenStudy (anonymous):

first substitute x and y values into the expression. 6(20)/3(2) - 4 now follow the order of operations: first do the multiplication 6(2)=12 3(2)=6 now do the division 12/6=2 now do the subtraction 2 - 4 = -2

OpenStudy (anonymous):

oops, soory don't look what i wrote i made a mistake!!

OpenStudy (anonymous):

6(20)=120 3(2)=6 120/6=20 20 - 4=16

OpenStudy (anonymous):

the answer is 16